Question 2 Two variables are defined, a regression equation is given, and one data point is given BAC-blood alcohol content( % of alcohol in the blood) Drinks number of alcoholic drinks BAC-0.01 24 + 0.015(Drinks) The data point is an individual who consumed 3 drinks and had a BAC of O.08. (a) Find the predicted value for the data point and compute the residual Enter the exact answers. Predicted value- Residual
Answer #1

Predicted value = -0.0124 + 0.015 * 3 = 0.0326

Residual = observed value - predicted value = 0.08 - 0.0326 = 0.0474